OpenStudy (anonymous):

let a denote the first term and r the common ratio of an exponential sequence. (i) Show that the general rule for finding the sum, Sn' of the fist n terms of the sequence is given by S_n=(a(1-r^n))/(1-r),r≠1. (ii) Hence deduce the sum to infinity of an exponential squence with |r|<1.

OpenStudy (atlas):

Write S_n and then find r*S_n See if you can see a pattern in them! You will be surprised

OpenStudy (anonymous):

Hmm, I don't understand

OpenStudy (atlas):

S_n = a + a*r + a*r^2 + a*r^3 + a*r^4.............nterms Do you understand this?

OpenStudy (anonymous):

yes, is that the starting

OpenStudy (atlas):

yeah.......now write r*S_n

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

rSn=ar+〖ar〗^2+〖ar〗^3+⋯+〖ar〗^(n-1)+〖ar〗^n

OpenStudy (atlas):

Ah great...........can you see most of the terms in S_n and r*S_n are same

OpenStudy (anonymous):

yes

OpenStudy (atlas):

great..now guess what will happen when you subtract them

OpenStudy (anonymous):

you mean Sn-rSn

OpenStudy (atlas):

yeah!!

OpenStudy (anonymous):

a-ar^n

OpenStudy (atlas):

So S -rS = a - ar^n right?? I am sure you can find S now

OpenStudy (anonymous):

good

OpenStudy (anonymous):

what of (ii)

OpenStudy (atlas):

oh it is very similar!! there will be infinite terms for both S and r*S.........so you need not care about the last terms

OpenStudy (atlas):

Try and write it out

OpenStudy (anonymous):

ok trying now

OpenStudy (anonymous):

hmm, no idea

OpenStudy (atlas):

what does infinity mean??

OpenStudy (anonymous):

no ending

OpenStudy (atlas):

yeah exactly .....Infinity means that 'continues for ever'

OpenStudy (atlas):

So if I write S for infinity terms S = a + ar + ar^2 + ar^3 ..............continues for ever

OpenStudy (atlas):

and then I can write rS = ar + ar^2 + ar^3 + .............continues for ever

OpenStudy (atlas):

Now what will happen if I subtract S and rS

OpenStudy (anonymous):

you are left with a

OpenStudy (atlas):

exactly so S-rS =a u can now find S for sure

OpenStudy (anonymous):

S is 1

OpenStudy (atlas):

oh no!!

OpenStudy (atlas):

S-rS =a so, S(1-r) =a (I have taken S common) Now find S

OpenStudy (anonymous):

S=a/(1-r)

OpenStudy (anonymous):

is this correct?

OpenStudy (atlas):

yeah right!!!

OpenStudy (anonymous):

Wow, atlas

OpenStudy (anonymous):

Thanks but I have questions

OpenStudy (anonymous):

more questions

OpenStudy (atlas):

yeah sure go on